. What is the pH of a solution of 6.5x10-1M HNO3? (5pts)
Answer #1

completely. HNO3 is a strong acid which dissociates HNO3 -> H+ + NO3- i [H+] = [HNO3] = 6·5x10-1M pt - log [H+] = -log(605x10Hope you will like my answer.